the 7th term of an AP is 19 and the 8th term is 22.find the sum of the first 40th terms of the series
Answer by Boreal(15235)

an=a1+(n-1)d
d=3, between 7th and 8th term
a7=a1+(n-1)d
a7=a1+6*3=a1+18
19=a1+18
a1=1
a40=1+(39)*3=118
sum=(40/2)(a1+a40)=20+(1+118)=119*20=2380